# Reasoning or Memorization? Unreliable Results of Reinforcement Learning Due to Data Contamination

### v2

@@ 4.2 Analysis of Memorization Capability
Meanwhile, we summarize the answer accuracy of the model under different ratios of prefix in Fig. 5. The Qwen2.5 models achieve remarkably high accuracy on MATH-500 even with partial questions. For instance, with 80% proportion of the questions, Qwen2.5-Math-7B reaches an accuracy of 63.8% on MATH-500. Even with only 40% proportion of the questions, the model still achieves an accuracy of 41.2%. Because our evaluation matches only the final numeric answer, the model can sometimes output the correct value by accident, which explains the anomalous accuracy of Llama on AIME2025 (solving exactly one problem with a faulty reasoning process). Besides, we also inspect several questions that Qwen solves correctly, as shown in Fig. 10 to 14 in Appendix. We find that the responses of Qwen contain coherent reasoning chains and even syntactically valid Python code, which, however, is not executed. The emergence of such structured solutions indicates that the training corpora may have included publicly available resources where benchmark problems are accompanied by detailed solutions. Besides, we provide more results for Qwen2.5 and Qwen3 under various sampling configurations in Appendix. The memorization of Qwen3 series is also analyzed in Appendix, showing similar results as Qwen2.5.
@@ 4.3 Spurious Rewards on RandomCalculation
Considering obtaining a reward of 1 on a RandomCalculation instance is virtually impossible, we report Max@16, the highest reward among 16 samples of initial model, in Fig. 7. We observe that on RandomCalculation datasets, Qwen2.5-Math-7B can surpass Max@16 when provided with correct reward signals. This finding indicates that reward-aligned RLVR effectively transfers the high-accuracy single-step arithmetic skills (as shown in Fig. 6) to more complex multi-step calculations, as illustrated by one reasoning trace in Fig. 15 of Appendix. In contrast, under the incorrect and random reward configurations, the Qwen model either maintains its base performance or exhibits only marginal and unstable improvements, due to the learning of format as explained in Shao et al. (2025). This improvements gradually disappear when calculation steps increased from 5 to 10. This further highlights the critical role of correct and well-aligned reward signals in enhancing model performance on uncontaminated and difficult datasets. Notably, Llama3.1-8B-Instruct fails to surpass the Max@16 even when trained with correct reward signals, and its accuracy falls below the greedy-decoding baseline when exposed to spurious signals. This discrepancy indirectly suggests that Qwen2.5 exhibits stronger mathematical capabilities compared to Llama3.1. However, such inherent strength is not the root cause of its performance boosting under spurious reward on contaminated datasets.
@@ 4.4 More Evidence for Memorization
Here, we provide more detailed analyses of Qwen’s sudden performance gains on MATH-500 under random reward. Let $\mathcal{J}_{\text{CLIP}}=\mathbb{E}_{\hat{A}_{i,t}}\!\Bigl{[}\min\left(r_{i,t}\hat{A}_{i,t},\text{clip}\left(r_{i,t},1-\epsilon,1+\epsilon\right)\hat{A}_{i,t}\right)\Bigr{]},$ where $\hat{A}_{i,t}$ is a random variable under the setup of random reward. Referring to Appendix B of  Shao et al. (2025), the gradient of the clipped policy has the following format:
@@
We further compare the responses of the model before and after RL, with Rouge-L and KL distance [^1] For each generated answer we compute $\mathrm{KL}\!\left(\mathrm{P}_{\text{Base}}\;\|\;\mathrm{P}_{\text{FT}}\right)$ over the full vocabulary, where $\mathrm{P}_{\text{Base}}$ and $\mathrm{P}_{\text{FT}}$ denote probability distributions produced by the fine-tuned and base models, respectively. as the similarity score. As shown in Tab. 3 and Fig. 8, the similarity in MATH-500 is substantially higher than in RandomCalculation, further implying that MATH-500 suffers from data contamination. Additionally, spurious rewards achieve even higher Rouge-L than correct reward after RL. Therefore, performance surge under spurious rewards arises because GRPO inadvertently triggers Qwen to retrieve memorized answers, rather than stimulating Qwen’s existing reasoning patterns like codes as explained in Shao et al. (2025). This is due to the exploitation bias of GRPO. However, RL with correct reward can still stimulate the model to find new reasoning paths, as verified with a smaller Rouge-L of response.
@@ Acknowledgments
Qin Liu is supported by the Amazon Nova Trusted AI Prize.
@@ Appendix E Quantitative Results for Qwen3
Appendix F Illustrative Examples of Model Output
Figures 10 and 11 present concrete instances of memorization by Qwen-2.5-7B on the AMC and AIME2024 benchmarks, respectively. Figures 1214 provide analogous examples for Qwen-2.5-Math-7B on MATH-500, AMC, and AIME2024. Figure 15 depicts a representative response produced by the RLVR-fine-tuned Qwen-2.5-Math-7B on the 5-step RandomCalculation benchmark.
@@ Appendix F Illustrative Examples of Model Output
[…]

### v3

@@ 4.2 Analysis of Memorization Capability
Meanwhile, we summarize the answer accuracy of the model under different ratios of prefix in Fig. 5. The Qwen2.5 models achieve remarkably high accuracy on MATH-500 even with partial questions. For instance, with 80% proportion of the questions, Qwen2.5-Math-7B reaches an accuracy of 63.8% on MATH-500. Even with only 40% proportion of the questions, the model still achieves an accuracy of 41.2%. Because our evaluation matches only the final numeric answer, the model can sometimes output the correct value by accident, which explains the anomalous accuracy of Llama on AIME2025 (solving exactly one problem with a faulty reasoning process). Besides, we also inspect several questions that Qwen solves correctly, as shown in Fig. 11 to 15 in Appendix. We find that the responses of Qwen contain coherent reasoning chains and even syntactically valid Python code, which, however, is not executed. The emergence of such structured solutions indicates that the training corpora may have included publicly available resources where benchmark problems are accompanied by detailed solutions. Besides, we provide more results for Qwen2.5 and Qwen3 under various sampling configurations in Appendix. The memorization of Qwen2.5 on LiveCodeBench (arxiv:LiveCodeBench) and Qwen3 series on math is also analyzed in Appendix, showing similar results as Tab. 2.
@@ 4.3 Spurious Rewards on RandomCalculation
Considering obtaining a reward of 1 on a RandomCalculation instance is virtually impossible, we report Max@16, the highest reward among 16 samples of initial model, in Fig. 7. We observe that on RandomCalculation datasets, Qwen2.5-Math-7B can surpass Max@16 when provided with correct reward signals. This finding indicates that reward-aligned RLVR effectively transfers the high-accuracy single-step arithmetic skills (as shown in Fig. 6) to more complex multi-step calculations, as illustrated by one reasoning trace in Fig. 16 of Appendix. In contrast, under the incorrect and random reward configurations, the Qwen model either maintains its base performance or exhibits only marginal and unstable improvements, due to the learning of format as explained in Spurious-Rewards. This improvements gradually disappear when calculation steps increased from 5 to 10. This further highlights the critical role of correct and well-aligned reward signals in enhancing model performance on uncontaminated and difficult datasets. Notably, Llama3.1-8B-Instruct fails to surpass the Max@16 even when trained with correct reward signals, and its accuracy falls below the greedy-decoding baseline when exposed to spurious signals. This discrepancy indirectly suggests that Qwen2.5 exhibits stronger mathematical capabilities than Llama3.1 before mid-training. However, such inherent strength is not the root cause of its performance boosting under spurious reward on contaminated datasets.
@@ 4.4 More Evidence for Memorization
Here, we provide more detailed analyses of Qwen’s sudden performance gains on MATH-500 under random reward. Let $\mathcal{J}_{\text{CLIP}}=\mathbb{E}_{\hat{A}_{i,t}}\!\Bigl[\min\left(r_{i,t}\hat{A}_{i,t},\text{clip}\left(r_{i,t},1-\epsilon,1+\epsilon\right)\hat{A}_{i,t}\right)\Bigr],$ where $\hat{A}_{i,t}$ is a random variable under the setup of random reward. Referring to Appendix B of  Spurious-Rewards, the gradient of the clipped policy has the following format:
@@
We further compare the responses of the model before and after RL, with Rouge-L and KL distance [^1] For each generated answer we compute $\mathrm{KL}\!\left(\mathrm{P}_{\text{Base}}\;\|\;\mathrm{P}_{\text{FT}}\right)$ over the full vocabulary, where $\mathrm{P}_{\text{Base}}$ and $\mathrm{P}_{\text{FT}}$ denote probability distributions produced by the fine-tuned and base models, respectively. as the similarity score. As shown in Tab. 3 and Fig. 8, the similarity in MATH-500 is substantially higher than in RandomCalculation, further implying that MATH-500 suffers from data contamination. Additionally, spurious rewards achieve even higher Rouge-L than correct reward after RL. Therefore, performance surge under spurious rewards arises because GRPO inadvertently triggers Qwen to retrieve memorized answers, rather than stimulating Qwen’s existing reasoning patterns like codes as explained in Spurious-Rewards. This is due to the exploitation bias of GRPO. However, RL with correct reward can still stimulate the model to find new reasoning paths, as verified with a smaller Rouge-L of response.
@@ Acknowledgments
The authors wish to thank the anonymous reviewers for their helpful comments. This work was partially funded by National Natural Science Foundation of China (No.62476061, 62376061, 62206057), Shanghai Rising-Star Program (23QA1400200), Natural Science Foundation of Shanghai (23ZR1403500). The computations were partially performed using Ascend AI Accelerators. The authors would like to thank Ascend Cloud Ecological Development Project for the support of Ascend 910 processors. Qin Liu is supported by the Amazon Nova Trusted AI Prize.
@@ Appendix E Quantitative Results for Qwen3
Appendix F RLVR on LiveMathBench
In the early exploratory phase, we conduct RLVR training on Qwen2.5-Math-7B with the clean LiveMathBench dataset. The experimental results are shown in Fig. 10. Under correct reward signal, the model achieves a limited improvement, which is mainly due to the relatively small amount of training data. In contrast, when using random reward signal, the model fails to obtain stable performance gains and ultimately exhibits a declining trend.
Appendix G Other Reasoning Domains
Beyond the mathematics domain, we also conduct preliminary memory tests on LiveCodeBench (arxiv:LiveCodeBench), a commonly used code evaluation benchmark. The results are shown in Tab. 14. When 80% proportion of problems provided as prompts, Qwen2.5-Math-7B is able to accurately reproduce 56.59% of the problems, whereas Llama3.1-8B only reproduces 4.40%. We also observe that the outputs of Qwen2.5-Math-7B frequently included complete test cases, as illustrated in Fig. 17.
Appendix H Illustrative Examples of Model Output
Figures 11 and 12 present concrete instances of memorization by Qwen2.5-7B on the AMC and AIME2024 benchmarks, respectively. Figures 1315 provide analogous examples for Qwen2.5-Math-7B on MATH-500, AMC, and AIME2024. Figure 16 depicts a representative response produced by the RLVR-fine-tuned Qwen2.5-Math-7B on the 5-step RandomCalculation benchmark.
Appendix I Hardware and Software Requirements
All experiments in this work were conducted on NVIDIA A800 80G GPUs, with Ubuntu 20.04.6 LTS as the operating system and CUDA driver version 12.4. Reinforcement learning (RL) training is performed using 8 A800 GPUs and 512 GB of RAM.
Appendix J RLVR Training Details
We follow the standard TTRL experimental configuration. Specifically, during training, the learning rate is set to 5e-7 and the temperature is set to 1.0. For each prompt, we generate 16 samples. The training batch size is set to 128. To improve efficiency and reduce memory usage, we enable FlashAttention to accelerate attention computation in the Transformer model.
@@ Appendix I Hardware and Software Requirements
Figure 10: RLVR training is performed on Qwen2.5-Math-7B with the LiveMathBench (v202412 + v202505) dataset under correct and random reward signals.
@@
AIME2025 30 55.44 13.33 36.99 0.00 32.92 0.00
MinervaMath 272 37.42 3.31 32.67 0.00 28.79 0.00
LiveMathBench 100 39.15 5.00 31.26 0.00 29.49 0.00
Table 9: Accuracy (Exact Match, EM) and ROUGE-L scores on several datasets (lower scores in gray) under different prompt prefix ratios in greedy decoding mode (Greedy (w/o Template) configuration).
@@
Model 80%-Problem 60%-Problem 40%-Problem
ROUGE-L EM ROUGE-L EM ROUGE-L EM
Qwen2.5-Math-7B 83.47 56.59 75.34 35.16 66.66 20.33
Qwen2.5-Math-7B-Instruct 46.60 4.95 36.49 0.00 32.62 0.00
Qwen2.5-7B 94.54 85.71 93.43 81.32 91.46 70.88
Qwen2.5-7B-Instruct 77.50 48.90 70.79 26.37 67.03 16.48
Llama3.1-8B 45.28 4.40 37.53 1.10 29.78 0.00
Llama3.1-8B-Instruct 36.20 1.65 32.77 0.00 28.33 0.00
Table 14: Accuracy (Exact Match, EM) and ROUGE-L scores of different models on LiveCodeBench (test generation) under different prompt prefix ratios in greedy decoding mode (Greedy (w/o Template) configuration).
Example from LiveCodeBench Problem: You are given a 0-indexed integer array nums. You have to find the maximum sum of a pair of numbers from nums such that the maximum digit in both numbers are equal. Return the maximum sum or -1 if no such pair exists. Test Cases: Input: [51, 71, 17, 24, 42] Output: 88 Input: [1, 2, 3, 4] Output: -1
Partial Prompt Prompt: You are given a 0-indexed integer array nums. You have to find the maximum sum of a pair of numbers from nums such that the maximum digit in both numbers are equal. Return the maximum
Qwen2.5-Math-7B Output sum or -1 if no such pair exists. Example 1:
Input: nums = [51,71,17,24,42] Output: 88
Explanation: For i = 1 and j = 2, nums[i] and nums[j] have equal maximum digits with a pair sum of 71 + 17 = 88. For i = 3 and j = 4, nums[i] and nums[j] have equal maximum digits with a pair sum of 24 + 42 = 66. It can be shown that there are no other pairs with equal maximum digits, so the answer is 88. Example 2:
Input: nums = [1,2,3,4] Output: -1
Explanation: No pair exists in nums with equal maximum digits. Constraints: 2 <= nums.length <= 100 1 <= nums[i] <= 104 Code 1class Solution { 2 public int maxSum(int[] nums) { 3 int max = -1; 4 for(int i=0;i<nums.length-1;i++){ 5 for(int j=i+1;j<nums.length;j++){ 6 if(maxDigit(nums[i])==maxDigit(nums[j])){ 7 max = Math.max(max,nums[i]+nums[j]); 8 } 9 } 10 } 11 return max; 12 } 13 public int maxDigit(int n){ 14 int max = 0; 15 while(n>0){ 16 int rem = n%10; 17 max = Math.max(max,rem); 18 n = n/10; 19 } 20 return max; 21 } 22}
Figure 17: Memorization Example of Qwen2.5-Math-7B on LiveCodeBench. In this case, the first 80% of the original problem is used as the prompt, and the generation is performed under the Greedy (w/o Template) configuration. Notably, the model even included the test cases in its output.